# Encoded probabilistic imaginary-time evolution on a trapped-ion quantum computer for ground and excited states of spin qubits

### v1

@@ III.2.3 Excited state of spin singlet
The total success probability for the PITE is shown in Fig. 8(b). The results of the noiseless simulation and encoded simulation (H1-1 and H1-1E) show similar values, but it was observed that the total success probability is smaller in the case without QED. The measured probability of Slater determinants is presented in Fig. 8(c). It can be confirmed that the classical fidelities of H1-1 and H1-1E using QED are 0.98 and 0.97, respectively, and that QED improves on the classical fidelity of 0.87 which does not use QED.
@@ IV Conclusions
In this study, the simulation results were discarded when quantum errors were detected using QED. However, this resulted in an increase in the variance of finite measurements. Therefore, it would be beneficial to implement quantum error correction. Furthermore, due to the constraints of the hardware, the minimum active space, comprising the $a$ and $e$ orbitals, was selected for the calculations. However, it would be more optimal to utilise a broader active space to enhance the precision of the calculations. As the number of orbitals included in the active space increases, the computational complexity rises exponentially, rendering classical computers inadequate for such calculations. Consequently, the potential of quantum computers is expected to expand considerably [120]. As PITE has already been demonstrated to be quadratically accelerated [66], it may be possible in the future to provide experimental evidence that quantum computation is more effective than classical computation in practical calculations based on advances in hardware.
@@ ACKNOWLEDGMENTS
The author acknowledges the contributions and discussions provided by the members of Quemix Inc. The authors thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was supported by JSPS KAKENHI under Grantin-Aid for Scientific Research No.21H04553, No.20H00340, and No.22H01517, JSPS KAKENHI. This work was partially supported by the Center of Innovations for Sustainable Quantum AI (JST Grant Number JPMJPF2221).
@@ C.1 Model parameters of the low-energy effective Hamiltonian
We summarized the model parameters of the low-energy effective Hamiltonian for $(\mathrm{Zr_{Al}V_{N}})^{0}$, $(\mathrm{Hf_{Al}V_{N}})^{0}$, $(\mathrm{Ti_{Al}V_{N}})^{0}$ complexes in w-AlN, and the NV--center in diamond in Table C. The model parameters were obtained using RESPACK [87]. The KS wave functions as an input of RESPACK were obtained by spin-restricted DFT calculations with the PBE and HSE functional starting from the optimized structure. We used cutoff energies for the dielectric functions of 3.0 Ry and 4.5 Ry for the NV center in diamond and complex defects in w-AlN, respectively. The energy bands are considered for 750 and 1300 bands for the NV--center in diamond and complex defects in w-AlN, respectively.
@@ C.2 Specific expression of Hamiltonian
$\Lambda_{2}=h_{ZZIZ}Z_{3}Z_{2}Z_{0}+h_{ZIIZ}Z_{3}Z_{0}+h_{IZZZ}Z_{2}Z_{1}Z_{0}+h_{ZZZZ}Z_{3}Z_{2}Z_{1}Z_{0}.$ (29)
@@
$V_{1}=h_{XXXZ}X_{3}X_{2}X_{1}Z_{0}+h_{ZZXZ}Z_{3}Z_{2}X_{1}Z_{0}+h_{ZXZZ}Z_{3}X_{2}Z_{1}Z_{0}+h_{XZZZ}X_{3}Z_{2}Z_{1}Z_{0}+h_{IXXZ}X_{2}X_{1}Z_{0}$
$+h_{ZXXZ}Z_{3}X_{2}X_{1}Z_{0}+h_{XZXZ}X_{3}Z_{2}X_{1}Z_{0}+h_{XIXZ}X_{3}X_{1}Z_{0}+h_{XXZZ}X_{3}X_{2}Z_{1}Z_{0}+h_{XXIZ}X_{3}X_{2}Z_{0}$
$+h_{IIXZ}X_{1}Z_{0}+h_{IXIZ}X_{2}Z_{0}+h_{XIIZ}X_{3}Z_{0}+h_{IXZZ}X_{2}Z_{1}Z_{0}+h_{ZXIZ}Z_{3}X_{2}Z_{0}$ (30)
@@
$V_{2}=h_{YYXZ}Y_{3}Y_{2}X_{1}Z_{0}+h_{YXYZ}Y_{3}X_{2}Y_{1}Z_{0}+h_{XYYZ}X_{3}Y_{2}Y_{1}Z_{0}+h_{IZXZ}Z_{2}X_{1}Z_{0}+h_{YZYZ}Y_{3}Z_{2}Y_{1}Z_{0}$
$+h_{YIYZ}Y_{3}Y_{1}Z_{0}+h_{YYZZ}Y_{3}Y_{2}Z_{1}Z_{0}+h_{YYIZ}Y_{3}Y_{2}Z_{0}+h_{ZIXZ}Z_{3}X_{1}Z_{0}+h_{XIZZ}X_{3}Z_{1}Z_{0}+h_{XZIZ}X_{3}Z_{2}Z_{0}.$ (31)
@@ D.2.1 Trunction error on qubit reduction using $\mathbb{Z}_{2}$ symmetry
$0.69\big{(}|e_{y\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle-|e_{y\downarrow}e_{x\uparrow}a_{1\uparrow}a_{1\downarrow}\rangle\big{)}+0.17\big{(}|e_{y\uparrow}e_{y\downarrow}e_{x\downarrow}a_{1\uparrow}\rangle-|e_{y\uparrow}e_{y\downarrow}e_{x\uparrow}a_{1\downarrow}\rangle\big{)}$ (32)
@@
$0.69\big{(}|e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle-|e_{y\uparrow}e_{y\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle\big{)}+0.17\big{(}|e_{y\downarrow}e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}\rangle-|e_{y\uparrow}e_{x\uparrow}e_{x\downarrow}a_{1\downarrow}\rangle\big{)}.$ (33)
@@ D.2.2 Ground-state calculation using emulator of H1-1
Table 7: Computed results for the ground state of the spin singlet state for NV center in diamond, TiV in w-AlN, and HfV in w-AlN on the emulator of H1-1 (H1-1E).
Total success probability Classical fidelity Discard rate

### v2

@@ III.2.3 Excited state of spin singlet
The total success probability for the PITE is shown in Fig. 8(b). The results of the noiseless simulation and encoded simulation (H1-1 and H1-1E) show similar values, but it was observed that the total success probability is smaller in the case without QED. The MSE is plotted in Fig. 8(c). Reflecting the small number of two qubit gates included in the circuit, the MSE for the result without QED is smaller than that for the ground state. The utilization of QED results in a reduction in MSE compared to instances where QED is not employed. The effectiveness of QED in the calculation of excited states was also observed.
The measured probability of Slater determinants and is MSE are presented in Fig. 8(d) and (e), respectively. It can be confirmed that the classical fidelities of H1-1 and H1-1E using QED are 0.98, and that QED improves on the classical fidelity of 0.87 which does not use QED.
@@ IV Conclusions
In this study, the simulation results were discarded when quantum errors were detected using QED. However, this resulted in an increase in the variance of finite measurements. Therefore, it would be beneficial to implement quantum error correction. We were unable to display the energy and fidelity in this study, which is an important quantity in the field of quantum chemistry, owing to the limited hardware resources. One calculation method with a constant depth overhad is to use quantum-state tomography, which causes scalability problems. Phase estimation enables the provision of energy with an additional circuit depth and avoids an increase in the number of measurements, which will be available in future protocols. In addition, since we successfully prepared the ground state, we will calculate such physical quantities, including one-body Green’s function and response function [27, 28]. Furthermore, due to the constraints of the hardware, the minimum active space, comprising the $a$ and $e$ orbitals, was selected for the calculations. However, it would be more optimal to utilize a broader active space to enhance the precision of the calculations. As the number of orbitals included in the active space increases, the computational complexity rises exponentially, rendering classical computers inadequate for such calculations. Consequently, the potential of quantum computers is expected to expand considerably [124]. As PITE has already been demonstrated to be quadratically accelerated [67], it may be possible in the future to provide experimental evidence that quantum computation is more effective than classical computation in practical calculations based on advances in hardware.
@@ ACKNOWLEDGMENTS
The author acknowledges the contributions and discussions provided by Prof. Shinji Tsuneyuki and the members of Quemix Inc. The authors thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was supported by JSPS KAKENHI under Grantin-Aid for Scientific Research No.21H04553, No.20H00340, and No.22H01517, JSPS KAKENHI. This work was partially supported by the Center of Innovations for Sustainable Quantum AI (JST Grant Number JPMJPF2221).
@@ C.1 Model parameters of the low-energy effective Hamiltonian
We summarized the model parameters of the low-energy effective Hamiltonian for $(\mathrm{Zr_{Al}V_{N}})^{0}$, $(\mathrm{Hf_{Al}V_{N}})^{0}$, $(\mathrm{Ti_{Al}V_{N}})^{0}$ complexes in w-AlN, and the NV--center in diamond in Table 5. The model parameters were obtained using RESPACK [88]. The KS wave functions as an input of RESPACK were obtained by spin-restricted DFT calculations with the PBE and HSE functional starting from the optimized structure. We used cutoff energies for the dielectric functions of 3.0 Ry and 4.5 Ry for the NV center in diamond and complex defects in w-AlN, respectively. The energy bands are considered for 750 and 1300 bands for the NV--center in diamond and complex defects in w-AlN, respectively.
@@ C.2 Specific expression of Hamiltonian
$\Lambda_{2}=h_{ZZIZ}Z_{3}Z_{2}Z_{0}+h_{ZIIZ}Z_{3}Z_{0}$
$+h_{IZZZ}Z_{2}Z_{1}Z_{0}+h_{ZZZZ}Z_{3}Z_{2}Z_{1}Z_{0}.$ (33)
@@
$V_{1}=h_{XXXZ}X_{3}X_{2}X_{1}Z_{0}+h_{ZZXZ}Z_{3}Z_{2}X_{1}Z_{0}$
$+h_{ZXZZ}Z_{3}X_{2}Z_{1}Z_{0}+h_{XZZZ}X_{3}Z_{2}Z_{1}Z_{0}$
$+h_{IXXZ}X_{2}X_{1}Z_{0}+h_{ZXXZ}Z_{3}X_{2}X_{1}Z_{0}$
$+h_{XZXZ}X_{3}Z_{2}X_{1}Z_{0}+h_{XIXZ}X_{3}X_{1}Z_{0}$
$+h_{XXZZ}X_{3}X_{2}Z_{1}Z_{0}+h_{XXIZ}X_{3}X_{2}Z_{0}+h_{IIXZ}X_{1}Z_{0}$
$+h_{IXIZ}X_{2}Z_{0}+h_{XIIZ}X_{3}Z_{0}+h_{IXZZ}X_{2}Z_{1}Z_{0}$
$+h_{ZXIZ}Z_{3}X_{2}Z_{0}$ (34)
@@
$V_{2}=h_{YYXZ}Y_{3}Y_{2}X_{1}Z_{0}+h_{YXYZ}Y_{3}X_{2}Y_{1}Z_{0}$
$+h_{XYYZ}X_{3}Y_{2}Y_{1}Z_{0}+h_{IZXZ}Z_{2}X_{1}Z_{0}+h_{YZYZ}Y_{3}Z_{2}Y_{1}Z_{0}$
$+h_{YIYZ}Y_{3}Y_{1}Z_{0}+h_{YYZZ}Y_{3}Y_{2}Z_{1}Z_{0}+h_{YYIZ}Y_{3}Y_{2}Z_{0}$
$+h_{ZIXZ}Z_{3}X_{1}Z_{0}+h_{XIZZ}X_{3}Z_{1}Z_{0}+h_{XZIZ}X_{3}Z_{2}Z_{0}.$ (35)
@@ D.2.1 Trunction error on qubit reduction using $\mathbb{Z}_{2}$ symmetry
$0.69\big{(}|e_{y\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle-|e_{y\downarrow}e_{x\uparrow}a_{1\uparrow}a_{1\downarrow}\rangle\big{)}$
$+0.17\big{(}|e_{y\uparrow}e_{y\downarrow}e_{x\downarrow}a_{1\uparrow}\rangle-|e_{y\uparrow}e_{y\downarrow}e_{x\uparrow}a_{1\downarrow}\rangle\big{)}$ (36)
@@
$0.69\big{(}|e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle-|e_{y\uparrow}e_{y\downarrow}a_{1\uparrow}a_{1\downarrow}\rangle\big{)}$
$+0.17\big{(}|e_{y\downarrow}e_{x\uparrow}e_{x\downarrow}a_{1\uparrow}\rangle-|e_{y\uparrow}e_{x\uparrow}e_{x\downarrow}a_{1\downarrow}\rangle\big{)}.$ (37)
@@ D.2.2 Ground-state calculation using emulator of H1-1
Table 7: Computed results for the ground state of the spin singlet state for NV center in diamond, TiV in w-AlN, and HfV in w-AlN on the emulator of H1-1 (H1-1E). Total success probability $\mathbb{P}_{\mathrm{tot}}$, classical fidelity $F$, and discard rate $d$ are summarized.
$\mathbb{P}_{\mathrm{tot}}$ $F$ $d$